# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ 1.3.2 Ordering and indices
For a $d$-dimensional vector $x\in\mathbb{R}^{d}$, we denote its coordinates as
@@
Similarly, $M[i,j]$ denotes the cell in the $i$th row and $j$th column of the matrix $M$. Furthermore, for multi-index $k=(k[1],k[2],\cdots,k[d])^{T}\in\mathbb{N}^{d}$, we use the convention $|k|=\sum_{j=1}^{d}k[j]$ and $k!=\prod_{j=1}^{d}k[j]!$.
For elements in $\mathbb{R}^{d}$, we define the lexicographical ordering $\prec$” as follow: for all $x=(x[1],x[2],\cdots,x[d])^{T}$ and $y=(y[1],y[2],\cdots,y[d])^{T}$, we say $x\prec y$ if either $\sum_{i}x[i]<\sum_{i}y[i]$ or $\sum_{i}x[i]=\sum_{i}y[i]$ and there exists a positive integer $d^{\prime}\leq d$, such that $x[d^{\prime}]<y[d^{\prime}]$ and $x[i]=y[i],\;\forall i\leq d^{\prime}-1$. Let us then denote by $k_{(i)}\in\mathbb{N}^{d}$ the $i$th multi-index in $\mathbb{N}^{d}$ with respect to the lexicographical ordering. Note that the first index is $k_{(1)}=(0,0,\cdots,0)^{T}$.
Finally, let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a $l$th order differentiable function. Then the derivative of $f$ with respect to multi-index $k\in\mathbb{N}^{d}$ at $x\in\mathbb{R}^{d}$ is denoted by
@@ 1.3.3 Sets, functions and functionals
For a finite set $A$, let $|A|$ denote its cardinality. For $t>0$, denote by $tA$ the set $\{tx:x\in A\}$. For a set $A=\{x_{1},x_{2},\cdots,x_{|A|}\}\subset\mathcal{X}$ and function $f:\mathcal{X}\to\mathbb{R}$, the vector version of $f$ is denoted as $\bm{f}(A)=(f(x_{1}),f(x_{2}),\cdots,f(x_{|A|}))^{T}\in\mathbb{R}^{|A|}.$
For a set $\Omega\subset\mathbb{R}^{d}$ and a function $f:\Omega\to\mathbb{R}$, let $\|f\|_{C^{\alpha}}$ denote the Hölder-$\alpha$ norm of $f$ and let $C^{\alpha}(\Omega)_{1}$ be the set of functions on $\Omega$ with Hölder-$\alpha$ norm bounded by $1$. Similarly, let $\|f\|_{W^{\alpha}}$ be the Sobolev-$\alpha$ norm (same as the $\|\cdot\|_{2,2,\alpha}$ norm in Definition C.6 of [22]) of $f$ and let $W^{\alpha}(\Omega)_{1}$ be the unit ball with respect to this norm. If $\Omega$ coincides with the domain $\mathcal{X}$ of the GP, we abbreviate the above two sets as $C^{\alpha}_{1}$ and $W^{\alpha}_{1}$.
For $\mathcal{X}_{n}\subset\mathbb{R}^{d}$ a finite set, the supremum and $\ell^{2}$ norm of a function $f:\mathbb{R}^{d}\to\mathbb{R}$ on $\mathcal{X}_{n}$ is denoted by $\|f\|_{\infty,n}=\sup_{x\in\mathcal{X}_{n}}|f(x)|$ and $\|f\|_{2,n}^{2}=\sum_{x\in\mathcal{X}_{n}}f(x)^{2}$, respectively. For two sets $\Omega_{1},\Omega_{2}$ , $p\in(0,+\infty]$ and an operator $M:\ell^{p}(\Omega_{1})\to\ell^{p}(\Omega_{2})$, the $\ell^{p}$ operator norm of $M$ is defined as $\|M\|_{p}=\sup_{f\in\ell^{p}(\Omega_{1})}\|M(f)\|_{p}/\|f\|_{p}$.
1.3.4 Other notations
We use $\iota$ to denote imaginary part of a complex number. For $\alpha\in\mathbb{R}$, let $\underline{\alpha}$ be the largest integer strictly smaller than $\alpha$. For two functions $f_{1}(n)$ and $f_{2}(n)$ of the sample size $n$, the notation $f_{1}(n)\lesssim f_{2}(n)$ means that there exists a universal constant $c>0$, such that $f_{1}(n)\leq cf_{2}(n)$. Let $\mathbbm{1}_{m}$ be a column vector with length $m$ and all elements equal to one. Finally, we denote by $Z_{1}\overset{d}{=}Z_{2}$ that the random variables $Z_{1},Z_{2}$ have the same distribution.
@@ 1.4 Organization of the paper
The rest of the paper is organized as follows. Section 2 introduces Vecchia approximations of GPs. Section 3 provides some background material about polynomials and introduces the concept of norming sets. Subsequently, Section 4 defines the Layered Norming DAG, providing a principled, theory driven choice of the underlying graphical structures. Section 5 presents some probabilistic properties for Matérn processes and the corresponding Vecchia Matérn processes. Building on these probabilistic results, in Section 6, we prove minimax contraction rates for appropriately scaled Vecchia GPs, both in the oracle and adaptive frameworks using hierarchical Bayesian approaches. Section 7 proceeds to demonstrate our theoretical findings with a numerical analysis on synthetic data sets. In Section 8, we discuss the implication of our results. All the proofs are deferred to the Supplementary Material.
@@ 2 Vecchia approximations of Gaussian Processes
First we describe the Gaussian Process regression model where we have carried out our analysis, and then introduce the Vecchia approximation of GPs. We consider Vecchia GPs both with fixed hyper-parameters and by endowing them with another layer of prior, resulting in a hierarchical prior distribution.
@@ 2.1 Gaussian Process Regression
Let us assume that we observe $Y_{1},...,Y_{n}$ satisfying
@@
where the covariates $(X_{i})_{1\leq i\leq n}$ are deterministic, belonging to a compact subset $\mathcal{X}\subset\mathbb{R}^{d}$, $f\in L_{2}(\mathcal{X})$ is the unknown functional parameter of interest, and $\sigma^{2}$ is the unknown variance of the noise. Let us denote by $\mathcal{D}_{n}=\{(X_{i},Y_{i})_{1\leq i\leq n}\}$ our data. We focus here on the Bayesian Gaussian process regression framework, by endowing $f$ with a GP prior, i.e.
@@
where $Z^{\tau,s}$ is a GP with rescaling parameters $(\tau,s)$. We also put a prior on the unknown parameter $\sigma^{2}$, denoted as
@@
The posterior distribution of the regression function given the data, i.e. $f|\mathcal{D}_{n}$, serves as a probabilistic solution to the regression problem. If the hyper-parameters $\tau,s$ and other tuning parameters of the Gaussian process are fixed, by conjugacy the posterior is also a Gaussian process. However, in practice, it is usually difficult to directly choose these hyper-parameters. Therefore, researchers often let the data decide about the optimal choice by endowing them with another layer of prior  [^2] There may exist other parameters for the process $Z^{\tau,s}_{x}$ other than $(\tau,s)$, for example smoothness $\alpha$ in the Matérn process in equation (2). However, fitting these hyper-parameters is often computationally more challenging., i.e.
@@
Equations (4) - (7) jointly specify the Gaussian process regression model. Due to the presence of the hyper-prior (7), the conjugacy is typically lost and the posterior is computed using Markov Chain Monte Carlo.
Even in the fixed hyper-parameter setting, where the GP posterior has an explicit form, the computation involves the inversion of an $n$ by $n$ matrix, which has computational complexity of order $O(n^{3})$. This complexity is intractable for many practical applications where $n$ can be as large as millions. Therefore, various approximation methods were proposed in the literature to speed up the computations. In the next section we describe one of the most popular such approach, the Vecchia GP.
@@ 2.2 Vecchia Gaussian Processes
Vecchia approximations of Gaussian Processes, or Vecchia GPs, are a type of methods that allow scalable computation for Gaussian processes. Specifically, a Vecchia GP $\hat{Z}^{\tau,s}$ is defined by two components, the mother Gaussian process $Z^{\tau,s}$ on $\mathcal{X}$ and a directed acyclic graph $G=(\mathcal{X},E)$. In most cases, $\mathcal{X}$ is an uncountable set and the directed acyclic graph $G$ can be split into two subgraphs: a graph on a finite reference set, and a bipartite graph between the reference set and the rest of the nodes. For simplicity, in this paper, we take the observed covariates $\mathcal{X}_{n}=(X_{1},...,X_{n})$ as the reference set. The graph on $\mathcal{X}_{n}$ characterizes the dependence structure of the reference set. The vertices of the bipartite graph on $\mathcal{X}$ consists of two parts: $\mathcal{X}_{n}$ and $\mathcal{X}\backslash\mathcal{X}_{n}$. All directed edges of this bipartite graph origin from a vertex in $\mathcal{X}_{n}$ and end at a vertex in $\mathcal{X}\backslash\mathcal{X}_{n}$. In other words, the nodes of $\mathcal{X}\backslash\mathcal{X}_{n}$ are conditional independent given $\mathcal{X}_{n}$ [^3] It is worth mentioning that the nodes of $\mathcal{X}\backslash\mathcal{X}_{n}$ are still marginally dependent. If the sample size $n$ is very large, the marginal dependency can be sufficiently strong to capture spatial patterns. .
The marginal densities of the Vecchia GP $\hat{Z}^{\tau,s}$ are defined via two steps. First, for the root vertex $X\in\mathcal{X}_{n}$, we have
@@
For all $X\in\mathcal{X}$, we say $X^{\prime}$ is a parent of $X$ if there is a directed edge from $X^{\prime}$ to $X$ in the DAG $G$. We denote the parent set of $X$, i.e., the collection of parents of $X$, as $\mathrm{pa}(X)$. Then $\hat{Z}^{\tau,s}_{X}$ given the process value at its parent locations follows the Gaussian distribution:
@@
Because $G$ is a directed acyclic graph, equations (8) and (2.2) consistently define the finite dimensional marginals of $\hat{Z}^{\tau,s}$, and hence define a valid Gaussian process, see for example Section S1 of [57] for a proof. Assuming $X_{1}$ is the root node, i.e., the node with no parent, we can proceed to write the joint density of the process $\hat{Z}^{\tau,s}$ on $\mathcal{X}_{n}$ as
@@
Intuitively speaking, the Vecchia approximated process $\hat{Z}^{\tau,s}$ is essentially obtained by removing some conditional dependencies in the mother process $Z^{\tau,s}$, where the DAG $G$ dictates which conditional dependencies (or equivalently, directed edges in $G$) to be kept. As a result, the precision matrices of the marginal distributions of $\hat{Z}^{\tau,s}$ and their Cholesky decompositions are sparse matrices. Such sparsity allows Vecchia GPs to perform scalable computation on large data sets. Specifically, if the cardinality of the parent sets $\mathrm{pa}(X_{i})$ for all covariates $X_{i}\in\mathcal{X}_{n}$ is bounded by $m$, then the computational complexity of evaluating the joint density on $\mathcal{X}_{n}$ via equation (12) is $O(nm^{3})$. This linear complexity in $n$ (for $m$ not increasing in $n$) is significantly smaller than the $O(n^{3})$ complexity of directly inverting the covariance matrix.
Endowing the regression function $f$ in equation (4) with the Vecchia GP
@@
the equations (4), (13) and (7) jointly specify the Vecchia GP regression model. Our paper focuses on analyzing the probabilistic and statistical properties of this model. It is evident that the statistical properties crucially depend on the parent sets $\mathrm{pa}(X_{i})$, or equivalently, the underlying DAG structure. While many different DAG structures are proposed based on various heuristics, we take a theory driven approach in selecting DAGs. Motivated by the flat limit of the conditional distribution of the GP with Matérn covariance kernel, we propose the Layered Norming DAG in the following two sections.
3 Polynomial interpolation and norming sets
The foundation of our probabilistic and statistical results is the until now unexplored connection of local polynomial interpolations and conditional distributions of Matérn processes. Therefore, we start by providing a brief introduction to polynomial interpolation. The proofs are deferred to Section A of the Supplementary Material. Several of the lemmas already appear in the literature, often without proofs. For completeness, we collect these results, refer to the monographs where they are stated, and provide aproof for them as well.
3.1 Vandermonde matrices and polynomial interpolation
We first introduce the multidimensional version of the Vandermonde matrices. For a finite subset $A=\{w_{1},w_{2},\cdots,w_{m}\}\subset\mathbb{R}^{d}$ and ordered multi-indices $k_{(1)},...,k_{(m)}\in\mathbb{N}^{d}$ (with respect to lexicographical ordering), the $d$-dimensional Vandermonde matrix is defined as
$V_{A}=\begin{bmatrix}1&1&\cdots&1\\ w_{1}^{k_{(2)}}&w_{2}^{k_{(2)}}&\cdots&w_{m}^{k_{(2)}}\\ w_{1}^{k_{(3)}}&w_{2}^{k_{(3)}}&\cdots&w_{m}^{k_{(3)}}\\ \vdots&\vdots&&\vdots\\ w_{1}^{k_{(m)}}&w_{2}^{k_{(m)}}&\cdots&w_{m}^{k_{(m)}}\end{bmatrix}$, (14)
where $w_{i}^{k_{(j)}}=\prod_{\ell=1}^{d}w_{i}[\ell]^{k_{(j)}[\ell]}$ and recalling that $k_{(1)}=(0,0,\cdots,0)^{T}$. Note that the standard, unidimensional Vandermonde matrix is a special case of the above $d$-dimensional version. Furthermore, for $x\in\mathbb{R}^{d}$, we define the vector $v_{x}$ as
$v_{x}=(1,{x}^{k_{(2)}},{x}^{k_{(3)}},\cdots,{x}^{k_{(m)}})^{T}.$ (15)
Next we discuss polynomial interpolation. Let $\mathscr{P}_{l}(\Omega)$ denote the space of (up to) $l$th order polynomials on $\Omega\subseteq\mathbb{R}^{d}$. Given $m$ interpolation points
$\{(w_{i},y_{i}):1\leq i\leq m,w_{i}\in\mathbb{R}^{d},y_{i}\in\mathbb{R}\}$
$P(w_{i})=y_{i},\forall 1\leq i\leq m.$
We are focusing here on “unisolvency”, when there exists a unique solution to the above problem. This topic has an over 100 years long history, see for instance the surveys [18] and [19]. The connection of unisolvency and Vandermonde polynomials has been stated in several sources, for example, in Proposition 3.5 of [54]. For completeness we provide a proof in the Supplementary Material.
Lemma 1.
Let $m={l+d\choose l}$, $A=\{w_{1},w_{2},\cdots,w_{m}\}\subset\mathbb{R}^{d}$ and $Y=(y_{1},y_{2},\cdots,y_{m})^{T}\in\mathbb{R}^{m}$. Then there exists a unique polynomial $P\in\mathscr{P}_{l}(\mathbb{R}^{d})$ satisfying $P(w_{i})=y_{i}$ if and only if the Vandermonde matrix $V_{A}$ is invertible. Moreover, this polynomial takes the form
$P(x)=Y^{T}V_{A}^{-1}v_{x}.$ (16)

### v4

@@ 1.3.2 Ordering and indices
For a $d$-dimensional vector $x\in\mathbb{R}^{d}$, we denote its coordinates by
@@
Similarly, $M[i,j]$ denotes the entry in the $i$th row and $j$th column of a matrix $M$. Furthermore, for multi-index $k=(k[1],k[2],\cdots,k[d])^{T}\in\mathbb{N}^{d}$, we use the conventions $|k|=\sum_{j=1}^{d}k[j]$ and $k!=\prod_{j=1}^{d}k[j]!$.
For vectors in $\mathbb{R}^{d}$, we define the lexicographical ordering $\prec$” as follows: for $x=(x[1],x[2],\cdots,x[d])^{T}$ and $y=(y[1],y[2],\cdots,y[d])^{T}$, we say that $x\prec y$ if either $\sum_{i}x[i]<\sum_{i}y[i]$ or $\sum_{i}x[i]=\sum_{i}y[i]$ and there exists $d^{\prime}\leq d$ such that $x[d^{\prime}]<y[d^{\prime}]$ and $x[i]=y[i],\;\forall i\leq d^{\prime}-1$. Let $k_{(i)}\in\mathbb{N}^{d}$ denote the $i$th multi-index in $\mathbb{N}^{d}$ with respect to this lexicographical ordering. Note that the first index is $k_{(1)}=(0,0,\cdots,0)^{T}$.
Finally, let $f:\mathbb{R}^{d}\to\mathbb{R}$ be an $l$th-order differentiable function. For any multi-index $k\in\mathbb{N}^{d}$, the corresponding partial derivative of $f$ at $x\in\mathbb{R}^{d}$ is denoted by
@@ 1.3.3 Sets, functions and functionals
For a finite set $A$, we denote its cardinality by $|A|$. For $t>0$, we write $tA=\{tx:x\in A\}$. For a set $A=\{x_{1},x_{2},\cdots,x_{|A|}\}\subset\mathcal{X}$ and a function $f:\mathcal{X}\to\mathbb{R}$, the vectorized version of $f$ over $A$ is $\bm{f}(A)=(f(x_{1}),f(x_{2}),\cdots,f(x_{|A|}))^{T}\in\mathbb{R}^{|A|}.$
For a set $\Omega\subset\mathbb{R}^{d}$ and a function $f:\Omega\to\mathbb{R}$, let $\|f\|_{C^{\alpha}}$ denote the Hölder-$\alpha$ norm of $f$ and let $C^{\alpha}(\Omega)_{1}$ be the unit ball in this norm. Similarly, let $\|f\|_{W^{\alpha}}$ be the Sobolev-$\alpha$ norm (equivalent to the $\|\cdot\|_{2,2,\alpha}$ norm in Definition C.6 of [22]) and denote by $W^{\alpha}(\Omega)_{1}$ the corresponding unit ball. When $\Omega$ coincides with the domain $\mathcal{X}$ of the GP, we abbreviate these sets by $C^{\alpha}_{1}$ and $W^{\alpha}_{1}$. The $L_{p}(\Omega)$ norm of the function $f$ is denoted by $\|f\|_{p}$.
For a finite set $\mathcal{X}_{n}\subset\mathbb{R}^{d}$, the supremum- and $L_{2}$-norm of a function $f:\mathbb{R}^{d}\to\mathbb{R}$ restricted to $\mathcal{X}_{n}$ are defined as $\|f\|_{\infty,n}=\sup_{x\in\mathcal{X}_{n}}|f(x)|$ and $\|f\|_{2,n}=[[|\mathcal{X}_{n}|^{-1}\sum_{x\in\mathcal{X}_{n}}f(x)^{2}]^{1/2}$, respectively. For sets $\Omega_{1},\Omega_{2}$ , $p\in(0,+\infty]$ and an operator $M:L_{p}(\Omega_{1})\to L_{p}(\Omega_{2})$, the $L_{p}$ operator norm is $\|M\|_{p}=\sup_{f\in L_{p}(\Omega_{1})}\|M(f)\|_{p}/\|f\|_{p}$.
1.3.4 Other notation
We use $\iota$ to denote imaginary unit. For $\alpha\in\mathbb{R}$, let $\underline{\alpha}$ denote the largest integer strictly smaller than $\alpha$. For two functions $f_{1}(n)$ and $f_{2}(n)$ of the sample size $n$, the notation $f_{1}(n)\lesssim f_{2}(n)$ means that there exists a universal constant $c>0$, such that $f_{1}(n)\leq cf_{2}(n)$ for all $n$. Let $\mathbbm{1}_{m}$ be the $m$-dimensional column vector whose entries are all equal to one. Finally, we denote by $Z_{1}\overset{d}{=}Z_{2}$ that the random variables $Z_{1},Z_{2}$ have the same distribution.
@@ 1.4 Organization of the paper
The rest of the paper is organized as follows. Section 2 introduces Vecchia approximations of GPs. Section 3 defines the layered norming DAG, providing a principled, theory driven choice of the underlying graphical structures. Section 4 establishes key probabilistic properties of Matérn processes and their Vecchia counterparts. Building on these results, Section 5 derives minimax contraction rates for appropriately rescaled Vecchia GPs, using both oracle tuning and hierarchical Bayesian approaches. Finally, Section 6 discusses the limitations and broader implications of our results. Due to space constraints, we defer the numerical analysis, some background material on polynomials and the concept of norming sets, as well as all proofs, to the Supplementary Material.
@@ 2 Vecchia approximations of Gaussian processes
We first describe the Gaussian Process regression model, and then introduce the Vecchia approximation for GPs. We consider Vecchia GPs both with fixed hyperparameters and with an additional layer of prior, resulting in a hierarchical prior distribution.
@@ 2.1 Gaussian process regression
Assume that we observe $Y_{1},...,Y_{n}$ satisfying
@@
where the covariates $(X_{i})_{1\leq i\leq n}$ are deterministic, and belong to a compact subset $\mathcal{X}\subset\mathbb{R}^{d}$, $f\in L_{2}(\mathcal{X})$ is the unknown functional parameter of interest, and $\sigma^{2}$ is the unknown noise variance. Let us denote by $\mathcal{D}_{n}=\{(X_{i},Y_{i})_{1\leq i\leq n}\}$ the observed data. We adopt a Bayesian Gaussian process regression framework by endowing $f$ with a GP prior, i.e.
@@
where $Z^{\tau,s}$ is a GP with rescaling parameters $(\tau,s)$. We also place a prior on the noise variance $\sigma^{2}$, denoted as
@@
The posterior distribution of the regression function given the data, $f|\mathcal{D}_{n}$, provides a probabilistic solution to the regression problem. If the hyperparameters $\tau,s$ and other GP tuning parameters are fixed, conjugacy ensures that the posterior is also a Gaussian process. In practice, however, choosing these hyperparameters directly is often difficult. Therefore, it is common to place another layer of prior on them, letting the data guide their selection  [^2] There may exist other hyperparameters for the process $Z^{\tau,s}_{x}$ beyond $(\tau,s)$, such as the regularity parameter $\alpha$ in the Matérn kernel (2). However, fitting these hyperparameters is often computationally more challenging., i.e.
@@
Equations (4) - (7) jointly define the Gaussian process regression model. Due to the presence of the hyper-prior (7), the conjugacy is typically lost and the posterior is computed via Markov Chain Monte Carlo.
Even in the fixed hyperparameter setting, where the GP posterior has an explicit form, the computation involves the inversion of an $n\times n$ matrix, which has computational complexity of order $O(n^{3})$. This complexity is intractable for many practical applications where $n$ can be as large as millions. Therefore, various approximation methods have been proposed in the literature to speed up the computations. In the next section, we describe one of the most popular such approaches: the Vecchia GP.
@@ 2.2 Vecchia Gaussian processes
Vecchia approximations of Gaussian processes, or Vecchia GPs, are a class of methods that allow scalable computation for Gaussian processes. Specifically, a Vecchia GP $\hat{Z}^{\tau,s}$ is defined by two components: the mother Gaussian process $Z^{\tau,s}$ on $\mathcal{X}$ and a directed acyclic graph $G=(\mathcal{X},E)$. In most cases, $\mathcal{X}$ is an uncountable set and the directed acyclic graph $G$ can be decomposed into two subgraphs: a graph defined on a finite reference set, and a bipartite graph between the reference set and the remaining nodes. For simplicity, in this paper we take the observed covariates $\mathcal{X}_{n}=(X_{1},...,X_{n})$ as the reference set. The graph on $\mathcal{X}_{n}$ characterizes the dependence structure within reference set. The vertices of the bipartite graph on $\mathcal{X}$ consist of two parts: $\mathcal{X}_{n}$ and $\mathcal{X}\backslash\mathcal{X}_{n}$. All directed edges in this bipartite graph originate from vertices in $\mathcal{X}_{n}$ and terminate at vertices in $\mathcal{X}\backslash\mathcal{X}_{n}$. In other words, the nodes of $\mathcal{X}\backslash\mathcal{X}_{n}$ are conditional independent given $\mathcal{X}_{n}$ [^3] It is worth noting that the nodes of $\mathcal{X}\backslash\mathcal{X}_{n}$ remain marginally dependent. When the sample size $n$ is large, this marginal dependency can be sufficiently strong to capture spatial patterns. .
The marginal densities of the Vecchia GP $\hat{Z}^{\tau,s}$ are defined in two steps. First, for the root vertex $X\in\mathcal{X}_{n}$, we have
@@
For any $X\in\mathcal{X}$, we say $X^{\prime}$ is a parent of $X$ if there is a directed edge from $X^{\prime}$ to $X$ in the DAG $G$. We denote the parent set of $X$, i.e., the collection of all parents of $X$, by $\mathrm{pa}(X)$. Then $\hat{Z}^{\tau,s}_{X}$ conditionally on the process value at its parent locations, follows the Gaussian distribution:
@@
Because $G$ is a directed acyclic graph, equations (8) and (2.2) consistently define the finite-dimensional marginals of $\hat{Z}^{\tau,s}$, and hence define a valid Gaussian process, see for example Section S1 of [62] for a proof. Assuming $X_{1}$ is the root node, i.e., the node with no parent, we can write the joint density of the process $\hat{Z}^{\tau,s}$ on $\mathcal{X}_{n}$ as
@@
Intuitively, the Vecchia approximated process $\hat{Z}^{\tau,s}$ is obtained by removing certain conditional dependencies in the mother process $Z^{\tau,s}$, where the DAG $G$ specifies which conditional dependencies (or equivalently, which directed edges) are retained. As a consequence, the precision matrices of the marginal distributions of $\hat{Z}^{\tau,s}$ and their Cholesky decompositions are sparse. This sparsity enables Vecchia GPs scalable computation on large data sets. Specifically, if the cardinality of the parent sets $\mathrm{pa}(X_{i})$ for all covariates $X_{i}\in\mathcal{X}_{n}$ is bounded by $m$, then the computational complexity of evaluating the joint density on $\mathcal{X}_{n}$ using equation (12) is $O(nm^{3})$. This linear complexity in $n$ (when $m$ does not increasing with $n$) is significantly smaller than the $O(n^{3})$ complexity of directly inverting the covariance matrix.
Endowing the regression function $f$ in equation (4) with the Vecchia GP prior
@@
the equations (4), (6), (13) and (7) together specify the Vecchia GP regression model. Our paper focuses on analyzing the probabilistic and statistical properties of this model. It is clear that these properties crucially depend on the underlying DAG structure. Although many different DAG structures have been proposed based on various heuristics, we take a theory driven approach to selecting DAGs. Motivated by the flat limit of the conditional distribution of Matérn processes, we propose the layered norming DAG in the following section.
3 Layered norming DAGs
The foundation of our probabilistic and statistical results is the previously underexplored connection between local polynomial interpolation and conditional distributions of Matérn processes. Due to space limitations, we present only the key concepts in this section, including norming sets and DAG structures. A brief introduction to this topic is provided in Section B of the supplement, where we define Vandermonde matrices and the unisolvency of polynomials, present useful lemmas, and discuss their connections. If the reader is not familiar with these concepts, we recommend consulting the supplement first.